OpenStudy (anonymous):

a rancher has 100 meters fo fencing to enclose two adjacent rectangular corrals ---approximate the dimensions such that the enclosed area will be 350 square meters (it also gives you the equations 4x+3y=100 and 2xy=A) i got as far as A=(200x-8x^2)/3 I don't know where to go from that

OpenStudy (dumbcow):

From there, substitute 350 for A (200x -8x^2)/3 = 350 Multiply by 3 on both sides 200x -8x^2 = 1050 set equal to 0 -8x^2 +200x -1050 = 0 Divide equation by 2 -4x^2 +100x -525 = 0 solve using quadratic formula

OpenStudy (anonymous):

okay I get 7.5 adn 17.5 for x now how do i find y?

OpenStudy (dumbcow):

plug those values back into one of the original equations --> 4(7.5) +3y = 100 --> 4(17.5) +3y = 100

OpenStudy (anonymous):

okay I got y= 10 and (70/3) but when you graph the equation you don't get the same answer how come?

OpenStudy (dumbcow):

oh because the graph is giving y-values that represent the Area, not the "y" dimensions

OpenStudy (anonymous):

oh okay thanks

OpenStudy (anonymous):

@marleythemongoose you forgot to multiply by 2

OpenStudy (anonymous):

I did. ;-) And was about to fix that. No need now. Well done.